# Logarithmic-Depth Quantum Circuits for Hamming Weight Projections

### v1

@@ I Introduction
Quantum computing is a computational paradigm that, for particular computational tasks, can potentially provide speedups to the best-known classical algorithms. Multiple classes of problems have now been shown to have quantum algorithms solving them faster than the current fastest classical algorithms. To leverage the power of quantum computing, several key properties, like entanglement and superposition, need to be manipulated effectively.
The Hamming distance of two binary strings is the number of substitutions needed to convert one string to the other. A special case of the Hamming distance is the Hamming weight, which is the Hamming distance of a string from the all-zeros string. The Hamming weight is more simply defined as the number of ‘1’ bits in the string, or, equivalently, the $l_{1}$-norm of a binary vector [1]. The Hamming weight of a string is a relevant quantity in multiple different fields such as cryptography [2], error correction [3], and information theory [4]. Many classical algorithms determine Hamming weight to facilitate applications such as internet-application lookup protocols [5] or determining the cryptographic performance of linear codes [6]. An entire suite of functions, called symmetric Boolean functions, rely solely on the Hamming weight of the input, encompassing foundational logical tasks such as AND, OR, Majority, and Parity [7]. Indeed, since classical computation primarily uses binary operations to perform calculations, the Hamming weight arises naturally in a variety of situations.
@@
The main contribution of our paper is a fast quantum algorithm for performing a coherent Hamming weight measurement, as a strongly related sequel to our Hamming-weight symmetry tests from [8, Sections 6.4.1 and 6.4.2]. [^1] Specifically, the following was noted after [8, Eq. (189)]: “As an aside, we note that a generalization of our method allows for performing a projection onto constant-Hamming-weight subspaces, which is useful in tasks like entanglement concentration.” “Coherent” in this context means that, not only is the classical measurement outcome equal to the Hamming weight, but it is also required for coherences in the same Hamming weight subspace to be preserved. In more detail, the aim of the coherent quantum Hamming weight measurement is to realize the following quantum channel on an $n$-qubit state $\rho$:
$\rho\to\sum_{x}|x\rangle\!\langle x|\otimes\Pi_{x}\rho\Pi_{x},$ (2)
@@
Prior work on the coherent Hamming weight measurement includes [10], which provides a linear-depth quantum circuit for this task. Interestingly, this circuit was considered experimentally in an ion-trap setup in [11]. Another work considered a coherent Hamming weight measurement as a collective magnetization measurement [12, 13], but there is no discussion of the circuit depth. In contrast to [10], our circuit has logarithmic depth and requires a logarithmic number of control qubits, assuming that qubit resets are allowed.
The ability to perform a coherent Hamming weight measurement has extensive applications in quantum information science. Realizing a coherent Hamming weight measurement efficiently is useful in information processing tasks like entanglement concentration [14] (see also [15, Chapter 19]), multiparty entanglement concentration [16], coherence concentration [17], multicasting of pure quantum states [18], and compression of pure states [19].
In this work, we present an algorithm that uses $n$ control qubits and has $\mathcal{O}(\log(n))$ depth to realize the coherent Hamming weight measurement on a quantum state of $n$ qubits. To keep the number of control qubits to be $n$, we assume the ability to perform qubit resets as the algorithm proceeds. Figure 7 depicts the corresponding quantum circuit. We also describe a trade-off between circuit depth and number of control qubits. The other end of the spectrum is an algorithm that uses $\mathcal{O}(\log(n))$ control qubits and has depth $\mathcal{O}(n\log(n))$.
Note on independent work—While finalizing our paper, we came across the concurrent and independent work in [20], which also provides shallow-depth circuits for computing Hamming weight coherently.
@@ II Preliminaries
To realize this measurement, we use a construction based on the idea of a pinching channel. See [21, Section 2.6.3] for a review of this concept, as well as the construction mentioned below. A pinching map is a channel of the form
@@
where $\{P_{x}\}_{x}$ for $x\in\{0,\ldots,N\}$ are orthogonal projectors that sum up to the identity. A pinching map has an alternative representation as a random unitary channel:
@@
$U_{y}\coloneqq\sum\limits_{x=0}^{N}\exp\!\left(\frac{2\pi\mathrm{i}yx}{N+1}\right)P_{x}$ (11)
and the equality results from basic Fourier analysis. The set $\{U_{y}\}_{y}$ is a unitary representation for the cyclic group $C_{N+1}$. To see this, we first define a cyclic group on $M$ elements:
$C_{M}=\langle a\mid a^{M}=e\rangle,$ (12)
@@
Thus, the set $\{U_{y}\}_{y}$ is a unitary representation for the group $C_{N+1}$.
The next order of business is to find a way to implement the set of unitaries $\{U_{y}\}_{y}$ on a quantum computer. For the case of Hamming-weight projections, each of the unitaries can be written as a product of $Z$ rotations on individual qubits, as shown in [8, Section 6.4.1]. Here we recall these points briefly. To this end, first consider the representation of the operator $R_{z}(\phi)$ in the computational basis:
@@
Thus, each unitary $U_{y}$ is composed of an overall phase and product of $Z$ rotations. The group structure of $\{U_{y}\}_{y}$ is more evident in this form, and the unitaries $U_{y}$ can be realized on a quantum computer.
@@
$=\alpha|00\rangle_{CS}+e^{\frac{-\mathrm{i}\phi}{2}}\beta|10\rangle_{CS}$ (28)
@@ III Algorithm Construction
III.1 First Algorithm with Zero-Padding
@@
Upon acting with the inverse quantum Fourier transform, the state becomes:
@@
Here, the group structure of $\{U_{y}\}_{y}$ is critical. We can replace this summation with a restricted sum over just the values of $y$ which are powers of two. The key insight is the fact that every $y$ in the sum can be written in terms of its binary representation. The group structure of the unitaries $\{U_{y}\}_{y}$ can then be leveraged to simplify the circuit. More concretely, we replace the summation with
@@
where $C_{m}$ denotes the $m$th qubit in the register $C$. This is a product of $\mathcal{O}(\log(n))$ elements, each controlled on a single qubit. In contrast, the general sum given above is a sum over $\mathcal{O}(n)$ elements, with multiple control qubits. To show the equivalence between the two operations, we consider the action of (44) on a state of the form $|a\rangle|\omega\rangle$, where $a$ is a bitstring.
@@
Now, expand $a$ in binary as $[a_{1}a_{2}\ldots a_{k}]$. We now notice that if any $a_{i}=1$, the modified Step 3 applies $U_{2^{i}}$. Thus, the overall effect is to apply the following unitary on the state $|\omega\rangle$:
$\prod_{\begin{subarray}{c}i=0\\ a_{i}=1\end{subarray}}^{k}U_{2^{i}}$ $=U_{\sum_{\begin{subarray}{c}i=0\\ a_{i}=1\end{subarray}}^{k}2^{i}}$ (50)
@@
where the first equality occurs because the unitaries all commute with each other, another property of the abelian group $C_{N+1}$. Thus, for each input bitstring $a$, the modified controlled-gates perform $U_{a}$ as needed.
@@
Figure 2: Practical implementation of the circuit from Figure 1. The individual phase angles are pulled out as rotations acting on the control qubits. In the figure $\alpha_{j}$ is a label for the rotation gate $R_{z}(\frac{\pi N}{N+1}2^{j-1})$ and $\beta_{j}$ is a label for $R_{z}(\frac{2\pi}{N+1}2^{j-1})$.
III.2 Second Algorithm without Zero-Padding
@@
Figure 3: Efficient practical implementation of the circuit from Figure 1. In the figure $\gamma_{j}$ is a label for the rotation gate $R_{z}(\frac{n\pi}{N+1}2^{j-1})$ and $\beta_{j}$ is a label for $R_{z}(\frac{2\pi}{N+1}2^{j-1})$.
@@
IV Depth-Width Tradeoff
@@
However, the tensor-product structure of $R_{z}^{\otimes n}(\beta_{j})$ allows for using an extra control qubits to reduce the depth. Once again, consider the first control qubit $C_{1}$. The control qubit $C_{1}$ can be encoded into a repetition code on $s$ qubits. More concretely, we tensor in $(s-1)$ qubits $A_{2},\ldots,A_{s}$ in the $|0\rangle$ state, and the control qubit $C_{1}$ is expanded to a control register $\bar{C}_{1}\equiv C_{1}A_{2}\ldots A_{s}$, as follows:
@@
This encoding can be achieved in constant quantum depth [22].
@@
To round out the algorithm, the $\operatorname{IQFT}$ and measurement need to be translated as well. Alternatively, we need to specify how to perform a repetition-code encoded $\operatorname{IQFT}$, measurement, and classical postprocessing to get the same measurement outcomes as the un-encoded algorithm. We now define the semi-classical encoded IQFT, making use of original insights from [25] (see also [26, Figure 1] for a visual depiction).
@@
[…]
@@
[…]
@@
[…]
@@
We now replace the Hadamard gates and measurement, with a tensor-product Hadamard operation, measurement, and classical post processing. Consider the action of a Hadamard gate and computational basis measurement on an arbitrary state. The probability of outcome $a\in\{0,1\}$ is given by
@@
$=\frac{1}{2}\left|\left(\langle 0|+(-1)^{a}\langle 1|\right)\left(\alpha|0\rangle+\beta|1\rangle\right)\right|^{2}$ (80)
@@
[…]
Proof. On each logical qubit, $F^{\text{sc}}_{k,s}$ consists of two operations, $H^{\otimes s}$ and a classically controlled rotation. In the traditional IQFT, the rotations are quantum-controlled and act on all qubits before, leading to an $\mathcal{O}(k^{2})$ depth. In the semi-classical version, we get a compiled rotation gate, leading to a linear depth.
The depth-optimal version of our algorithm for an $n$-qubit input state is shown in Figure 6. An important point to note here is that, even though Figure 6 depicts $nk$ control qubits, if qubit resets are available, we only need $n$ control qubits. This is because each step can be performed in a sequential manner as follows and in constant quantum depth (see Figure 7 for this variation): first $n$ qubits are prepared in a GHZ state, then the controlled rotations labeled by $\beta_{1}$ are performed, then the first step of the semiclassical encoded IQFT is performed, leading to a measurement outcome $a_{1}$. Then the $n$ control qubits are reset, prepared in a GHZ state, the controlled rotations labeled by $\beta_{2}$ are perform, then the second step of the semiclassical encoded IQFT is performed (with an action conditioned on $a_{1}$). There are $k-2$ more steps like this, thus leading to a $\mathcal{O}(\log n)$ depth circuit with $n$ control qubits.
@@ V Resource Counting
Table 1: Resource count for the algorithm construction for varying $s$ value that characterizes the depth-width tradeoff. We note that for the different estimates, we assume that qubit resets are available.

### v2

@@ I Introduction
Quantum computing is a computational paradigm that, for particular computational tasks, can potentially provide speedups to the best-known classical algorithms. Multiple classes of problems have now been shown to have quantum algorithms solving them faster than the current fastest classical algorithms [1]. To leverage the power of quantum computing, several key properties, like entanglement and superposition, need to be manipulated effectively.
The Hamming distance of two binary strings is the number of substitutions needed to convert one string to the other. A special case of the Hamming distance is the Hamming weight, which is the Hamming distance of a string from the all-zeros string. The Hamming weight is more simply defined as the number of ‘1’ bits in the string, or, equivalently, the $\ell_{1}$-norm of a binary vector [2]. The Hamming weight of a string is a relevant quantity in multiple different fields such as cryptography [3], error correction [4], and information theory [5]. Many classical algorithms determine Hamming weight to facilitate applications such as internet-application lookup protocols [6] or determining the cryptographic performance of linear codes [7]. An entire suite of functions, called symmetric Boolean functions, rely solely on the Hamming weight of the input, encompassing foundational logical tasks such as AND, OR, Majority, and Parity [8]. Indeed, since classical computation primarily uses binary operations to perform calculations, the Hamming weight arises naturally in a variety of situations.
@@
The main contribution of our paper is a fast quantum algorithm for performing a coherent Hamming weight measurement, as a strongly related sequel to our Hamming-weight symmetry tests from [9, Sections 6.4.1 and 6.4.2]. [^1] Specifically, the following was noted after [9, Eq. (189)]: “As an aside, we note that a generalization of our method allows for performing a projection onto constant-Hamming-weight subspaces, which is useful in tasks like entanglement concentration.” “Coherent” in this context means that, not only is the classical measurement outcome equal to the Hamming weight, but it is also required for coherences in the same Hamming weight subspace to be preserved. In more detail, the aim of the coherent quantum Hamming weight measurement is to realize the following quantum channel on an $n$-qubit state $\rho$:
$\rho\mapsto\sum_{x}|x\rangle\!\langle x|\otimes\Pi_{x}\rho\Pi_{x},$ (2)
@@
Prior work on the coherent Hamming weight measurement includes [11], which provides a linear-depth quantum circuit for this task. Interestingly, this circuit was considered experimentally in an ion-trap setup in [12]. Another work considered a coherent Hamming weight measurement as a collective magnetization measurement [13, 14], but there is no discussion of the circuit depth. In contrast to [11], our circuit has logarithmic depth and requires a linear number of control qubits, assuming that qubit resets are allowed.
One of the key components of our algorithm, that allows for the reduction in depth, is the quantum fan-out operation, implemented in constant quantum-depth using mid-circuit measurements. The usage of the quantum fan-out is known to help in various other algorithms [15], including factoring using Shor’s algorithm [16]. Another component of our algorithm involves using controlled rotation gates to calculate Hamming weights coherently, which bears similarities to the earlier work of [17].
The ability to perform a coherent Hamming weight measurement has various applications in quantum information science. Realizing a coherent Hamming weight measurement efficiently is useful in information processing tasks like entanglement concentration [18] (see also [19, Chapter 19]), multiparty entanglement concentration [20], coherence concentration [21], multicasting of pure quantum states [22], and compression of pure states [23]. Another application of the coherent Hamming weight measurement is in error detection and correction. In the natural setting of independent and identically distributed (i.i.d.) amplitude damping noise, a coherent Hamming weight measurement can used to detect which error has occurred when using the constant-excitation codes of [24]. For error correction of i.i.d. amplitude damping noise, when using such codes, it is possible to perform the correction step based on the constant-excitation space projected into, by employing a coherent Hamming weight measurement.
In this work, we present an algorithm that uses $n$ control qubits and has $\mathcal{O}(\log(n))$ depth to realize the coherent Hamming weight measurement on a quantum state of $n$ qubits. To keep the number of control qubits to be $n$, we assume the ability to perform mid-circuit measurements and qubit resets as the algorithm proceeds. Dynamic circuits on IBM hardware were used in [25] and compiling quantum circuits to use qubit resets was employed in [26].
Figure 7 depicts the corresponding quantum circuit. We also describe a trade-off between circuit depth and number of control qubits. The other end of the spectrum is an algorithm that uses $\mathcal{O}(\log(n))$ control qubits and has depth $\mathcal{O}(n\log(n))$.
Note on independent work—While finalizing our paper, we came across the concurrent and independent works [27, 28], which also provide shallow-depth circuits for computing Hamming weight coherently.
@@ II Preliminaries
To realize this measurement, we use a construction based on the idea of a pinching channel. See [29, Section 2.6.3] for a review of this concept, as well as the construction mentioned below. A pinching map is a channel of the form
@@
where $\{P_{x}\}_{x}$ for $x\in\{0,\ldots,N\}$ are orthogonal projectors that sum up to the identity. Importantly, a pinching map has an alternative representation as a random unitary channel:
@@
$U_{y}\coloneqq\sum\limits_{x=0}^{N}\exp\!\left(\frac{2\pi\mathrm{i}yx}{N+1}\right)P_{x},$ (11)
and the equality results from basic Fourier analysis. The set $\{U_{y}\}_{y}$ is a unitary representation of the cyclic group $C_{N+1}$. To see this, we first define a cyclic group on $M$ elements:
$C_{M}\coloneqq\langle a\mid a^{M}=e\rangle,$ (12)
@@
Thus, the set $\{U_{y}\}_{y}$ is a unitary representation of the group $C_{N+1}$.
The next order of business is to find a way to implement the set of unitaries $\{U_{y}\}_{y}$ on a quantum computer. For the case of Hamming-weight projections, each of the unitaries can be written as a product of $Z$ rotations on individual qubits, as shown in [9, Section 6.4.1]. Here we recall these points briefly. To this end, first consider the representation of the operator $R_{z}(\phi)$ in the computational basis:
@@
Thus, each unitary $U_{y}$ is composed of an overall phase and a product of $Z$ rotations. The group structure of $\{U_{y}\}_{y}$ is more evident in this form, and the unitaries $U_{y}$ can be realized on a quantum computer.
@@ III Algorithm Construction
The first algorithm uses zero padding to have exactly $2^{k}$ measurement outcomes. This simplifies the analysis and the proof, but it requires more ancilla qubits and rotation gates. The second algorithm eliminates the extra zero-padded ancilla qubits and rotation gates. The analysis of this algorithm builds upon the analysis of the first algorithm.
III.1 First Algorithm with Zero Padding
@@
Upon acting on the control qubits with the inverse quantum Fourier transform, the state becomes:
@@
Here, the group structure of $\{U_{y}\}_{y}$ is critical. We can replace this summation with a restricted sum over just the values of $y$ that are powers of two. The key insight is the fact that every $y$ in the sum can be written in terms of its binary representation. The group structure of the unitaries $\{U_{y}\}_{y}$ can then be leveraged to simplify the circuit. More concretely, we replace the summation with
@@
where $C_{m}$ denotes the $m$th qubit in register $C$. This is a product of $\mathcal{O}(\log(n))$ elements, each controlled on a single qubit. In contrast, the general sum given above is a sum over $\mathcal{O}(n)$ elements, with multiple control qubits. To show the equivalence between the two operations, we consider the action of (43) on a state of the form $|a\rangle|\omega\rangle$, where $a$ is a bitstring.
@@
Now, expand $a$ in binary as $[a_{1}a_{2}\ldots a_{k}]$. We notice that if any $a_{i}=1$, the modified Step 3 applies $U_{2^{i}}$. Thus, the overall effect is to apply the following unitary on the state $|\omega\rangle$:
$\prod_{\begin{subarray}{c}i=0\\ a_{i}=1\end{subarray}}^{k}U_{2^{i}}$ $=U_{\sum_{\begin{subarray}{c}i=0\\ a_{i}=1\end{subarray}}^{k}2^{i}}$
@@
where the first equality holds because the unitaries all commute with each other, another property of the abelian group $C_{N+1}$. Thus, for each input bitstring $a$, the modified controlled-gates perform $U_{a}$ as needed.
@@
Figure 2: Practical implementation of the circuit from Figure 1. The individual phase angles are pulled out as rotations acting on the control qubits. In the figure, $\alpha_{j}$ is a label for the rotation gate $R_{z}(\frac{\pi N}{N+1}2^{j-1})$ and $\beta_{j}$ is a label for $R_{z}(\frac{2\pi}{N+1}2^{j-1})$.
III.2 Second Algorithm without Zero Padding
@@
Figure 3: Efficient practical implementation of the circuit from Figure 1. In the figure, $\gamma_{j}$ is a label for the rotation gate $R_{z}(\frac{n\pi}{N+1}2^{j-1})$ and $\beta_{j}$ is a label for $R_{z}(\frac{2\pi}{N+1}2^{j-1})$.
@@
IV Depth-Width Tradeoff and 2D Architecture Implementation
@@
However, the tensor-product structure of $R_{z}^{\otimes n}(\beta_{j})$ allows for using extra control qubits to reduce the depth. Once again, consider the first control qubit $C_{1}$. The control qubit $C_{1}$ can be encoded into a repetition code on $s$ qubits. More concretely, we tensor in $(s-1)$ qubits $A_{2},\ldots,A_{s}$ in the $|0\rangle$ state, and the control qubit $C_{1}$ is expanded to a control register $\bar{C}_{1}\equiv C_{1}A_{2}\ldots A_{s}$, as follows:
@@
This encoding can be achieved in constant quantum depth by allowing for mid-circuit measurements and feedback [30].
@@
To round out the algorithm, the $\operatorname{IQFT}$ and measurement need to be translated as well. Alternatively, we need to specify how to perform a repetition-code encoded $\operatorname{IQFT}$, measurement, and classical postprocessing to get the same measurement outcomes as the un-encoded algorithm. We now define the semi-classical encoded IQFT, making use of original insights from [33] (see also [34, Figure 1] for a visual depiction).
@@
Figure 8: $2$-D implementation of the circuit given in Figure 7. We show the diagram for a $16$-qubit input state $|\psi\rangle$. Furthermore, we set $s=16$ to get the depth-optimal version of the algorithm. In the different operations, the label $\bar{a}$ refers classical controlled $Z$-rotation $R_{z}\!\left(\frac{\pi}{2^{k-1}}\bar{a}\right)$ where $\bar{a}\coloneqq a_{k-1}\cdots a_{1}$. Each bit $a_{i}$ is the sum of entries of the bitstring $b^{i}$ modulo $2$.
@@
The semi-classical encoded IQFT $F^{\text{sc}}_{k,s}$ acts on a $k$-qubit quantum state encoded into an $s$-qubit repetition code and has a recursive circuit description consisting of the following steps:
@@
    Apply $F^{\text{sc}}_{k-1,s}$ on the first $(k-1)s$ qubits, and obtain the measurement outcome $a_{1}\cdots a_{k-1}$.
@@
The semi-classical encoded IQFT $F^{\text{sc}}_{k,s}$ acting on an encoded state gives the same measurement outcome probabilities as performing the IQFT on the original quantum state and measuring in the computational basis.
@@
We now replace the Hadamard gates and measurement with a tensor-product Hadamard operation, measurement, and classical post processing. Consider the action of a Hadamard gate and a computational basis measurement on an arbitrary state. The probability of outcome $a\in\{0,1\}$ is given by
@@
The depth of the semi-classical encoded IQFT $F^{\text{sc}}_{k,s}$ is linear in $k$ and independent of $s$.
Proof. On each logical qubit, $F^{\text{sc}}_{k,s}$ consists of two operations: $H^{\otimes s}$ and a classically controlled rotation. In the traditional IQFT, the rotations are quantum-controlled and act on all qubits before, leading to an $\mathcal{O}(k^{2})$ depth. In the semi-classical version, we get a compiled rotation gate, leading to a linear depth.
The depth-optimal version of our algorithm for an $n$-qubit input state is shown in Figure 6. An important point to note here is that, even though Figure 6 depicts $nk$ control qubits, if qubit resets are available, we only need $n$ control qubits. This is because each step can be performed in a sequential manner as follows and in constant quantum depth (see Figure 7 for this variation): the first $n$ qubits are prepared in a GHZ state, then the controlled rotations labeled by $\beta_{1}$ are performed, then the first step of the semiclassical encoded IQFT is performed, leading to a measurement outcome $a_{1}$. Then the $n$ control qubits are reset, prepared in a GHZ state, the controlled rotations labeled by $\beta_{2}$ are performed, then the second step of the semiclassical encoded IQFT is performed (with an action conditioned on $a_{1}$). There are $k-2$ more steps like this, thus leading to a $\mathcal{O}(\log n)$ depth circuit with $n$ control qubits.
An implementation of the depth-optimal version on a 2D architecture is given in Figure 8. We note that the overall growth of the depth is still $\mathcal{O}(\log n)$ when restricted to a 2D ladder architecture. One leg of the ladder consists of ancilla qubits, and the other leg consists of the data qubits.
@@ V Resource Counting
Table 1: Resource count for the algorithm construction for varying $s$ values, the latter of which characterizes the depth-width tradeoff. We note that for the different estimates, we assume that qubit resets are available.